OpenStudy (anonymous):

(Geometry) Please put the equation in standard form. Containing E(4, 3) and F(6, 1)

OpenStudy (zzr0ck3r):

y = mx+b 3 = m4 + b b = 3 - m4 1 = m6+3-m4 1 = 2m + 3 m = -1 b = 7 y = -x+7

OpenStudy (zzr0ck3r):

make since?

OpenStudy (anonymous):

Got it, Thanks!